please help
4/5x + 2 greater or equal to 2/5
and
1-6x less than or equal to 7

Answer by ankor@dixie-net.com(22740)

please help
4/5x + 2 greater or equal to 2/5
:
Write it:
(4/5)x + 2 >= 2/5
:
Get rid of those annoying denominators, multiply equation by 5 and you have:
4x + 10 >= 2
:
Subtract 10 from both sides:
4x >= 2 -10
:
4x >= -8
:
Divide both sides by 4
x >= -8/4
:
x >= -2
:
:
1-6x less than or equal to 7
1 - 6x <= 7
:
Subtract 1 from both sides and you have:
-6x <= 7 - 1
:
-6x <= 6
:
We must have the variable (x) to be positive, so we multiply equation by -1
When we do this, we must reverse the inequality sign from < to >
+6x >= -6
:
Divide both sides by 6
x >= -6/6
:
x >= -1
:
What about this? Did this make sense to you now?